OpenStudy (anonymous):

Given 9801 = 3x3x3x3x11x11 , find √9801 (square root i think)

OpenStudy (mimi_x3):

99

OpenStudy (anonymous):

how did you work that out ?

OpenStudy (mimi_x3):

I used a calculator xD

OpenStudy (anonymous):

ah , what did you press ?

OpenStudy (mimi_x3):

LOL, pressed on squareroot sign then pressed the numbers, common sense anyway (:

OpenStudy (mimi_x3):

The answer is 99

OpenStudy (mimi_x3):

not 14 xD

OpenStudy (anonymous):

yeah that is what i mean , geez

OpenStudy (anonymous):

Looking at the 3x3x3x3x11x11, you can see that there are two pairs of 3s and one pair of 11s, so you could rewrite that as:\[3^{2} \times 3^{2} \times 11^{2}\]So you can see that multiplying those three numbers together (3, 3, and 11) will give you your answer:\[3 \times 3 \times 11 = 99\]

OpenStudy (anonymous):

ooohh thank you so much alex !!!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

its 99

OpenStudy (anonymous):

so with 2x2x7x7 would it basically be 2 x 7 = 14 and that is the square root of 196 ??

OpenStudy (anonymous):

Right, and if you were looking for the cubed root of something, you would need to look for triplets of its prime-number factors

OpenStudy (anonymous):

ok i have this other question which says : write 1764 as the product of its prime factors and then find the square root of 1764 ..

OpenStudy (anonymous):

1764 is divisible by 2 (a prime number), so add that to the list. Dividing it by 2 gives you 882, which is divisible by 2 also, so add 2 to the list again. 882/2=441, which is not divisible by 2 so you move on to the next prime number, 3. 441 is divisible by 3, so you add that to the list, and when divided gives you 147 which is divisible by 3 also so you add 3 again. You do this continually until you get 1. \[2 \times 2 \times 3 \times 3 \times 7 \times 7 = 1764\] So \[2^{2} \times 3^{2} \times 7^{2} = 1764\]And\[2 \times 3 \times 7 = \sqrt{1764} = 42\]

OpenStudy (anonymous):

thank you so much that is all from me today